OpenStudy (devonhoward15):

Drag a statement or reason to each box to complete the proof. If 4x+5=174x+5=17, then x=3x=3.

OpenStudy (devonhoward15):

@jhonyy9

jhonyy9 (jhonyy9):

4x+5=17 subtract from both sides 5 so what will get ?

OpenStudy (devonhoward15):

I think the first one is 4x+5=17 2-subtraction property of equality 3,x=3 4-bottom division one 5 idk

OpenStudy (devonhoward15):

that look right?

jhonyy9 (jhonyy9):

4x+5=17 subtract 5 from both sides and will get 4x=12 divide both sides by 4 and will get x = 3 hope helped

OpenStudy (devonhoward15):

@jhonyy9 please check

jhonyy9 (jhonyy9):

like how i ve wrote it above sure good luck